OpenStudy (vespersaviour):

A cylinder has a radius of 10 cm and a height of 8 cm. What is the volume of the cylinder?

OpenStudy (stbest):

2513.27 because 2 find the volume you need 2 do πr2h

OpenStudy (vespersaviour):

\[A. 160\pi cm ^{3}\] \[B. 480\pi cm ^{3}\] \[C.640\pi cm ^{3}\] \[D.800\pi cm ^{3}\] ^^the answer choices

OpenStudy (vespersaviour):

@563blackghost can you help? i think its D

563blackghost (563blackghost):

You are correct ^^

OpenStudy (vespersaviour):

^.^ tank chu

563blackghost (563blackghost):

np :)